Prove that the planes

7x+ 4y− 4z+ 30 = 0,36x−51y+12z+17 = 0,14x+8y−8z−12 = 0 and 12x−17y+4z−3 = 0 form the four faces of a cuboid.


please solve this ASAP.


1
Expert's answer
2021-06-08T18:32:48-0400

Given


  1. 7x + 4y - 4z =-30
  2. 14x + 8y -8z = 12 <=> 7x + 4y - 4z = 6
  3. 36x - 51y + 12z = -17 <=> 12x - 17y +4z = -17/3
  4. 12x -17y + 4z = 3


Both planes are parallel if condition suffices



"\frac{a_1}{a_2} = \frac{b_1}{b_2} = \frac{c_1}{c_2}."


where a1x + b1y + c1z = dand a2x + b2y + c2z = dare planes.


Let's consider 1st and 2nd equation.



"\frac{7}{7} = \frac{4}{4} = \frac{-4}{-4}"

Let's consider 3th and 4th equation.



"\frac{12}{12} = \frac{-17}{-17} = \frac{4}{4}"

So, 1,2 and 3,4 planes are parallel. A cuboid is a figure bounded by six face parallel to each opposite one. So, the planes form cuboid.